# A Greedy Quantum Route-Generation Algorithm

### v1

@@ VII QPU Results
This work presents the first effective algorithm for generating routes using an annealing-based approach. We demonstrated its efficiency by solving the FSVRPTW. We prove that it converges to a feasible solution, and we benchmark it against state-of-the-art classical and hybrid annealing-based approaches using D-Wave. We showed it enjoys a smaller optimality gap than other annealing-based approaches, at a competitive computation time, even for practical-sized problems of 50 customers. It also shows to be noise-robust when executed on a QPU, by taking advantage of the entire sampleset returned by the annealer. Natural directions for future work include determining the dependence of the proposed algorithm’s performance on the dataset, e.g. using different time discretizations and graphs generated from different distributions. We may also explore the effect of the control parameters, e.g. $\theta$. These future directions will provide insight into how this adaptive algorithm may be used to solve a more general class of problems outside of routing and scheduling.
VIII Acknowledgements
J.M. acknowledges Ketan Savla, and David E. Bernal Neira for helpful input and discussion during the development of these ideas.

### v2

@@ 7 Conclusions and Future Work
This work presents the first effective algorithm for generating routes using an annealing-based approach. We demonstrated its efficiency by solving the Fleet Sizing Vehicle Routing Problem with Time Windows (FSVRPTW). We prove that it converges to a feasible solution, and we benchmark it against state-of-the-art classical and hybrid annealing-based approaches. We showed that it enjoys a smaller optimality gap than other annealing-based approaches at a competitive computation time, even for practical-sized problems. It also shows to be noise-robust when executed on a QPU, taking advantage of the entire sample set returned by the annealer. Natural directions for future work include determining the dependence of the proposed algorithm’s performance on the dataset, e.g., using different time discretizations and graphs generated from other distributions. We may also explore the effect of the control parameters, e.g., $\theta$. These future directions will provide insight into how this adaptive algorithm may be used to solve a more general class of problems outside of routing and scheduling.
8 Acknowledgements
This work was supported in part by the PWICE fellowship of USC.